OpenStudy (anonymous):

HELP PLEASE! WILL MEDAL!!! I promise this will only take second =) The function H(t) = -16t2 + vt + s shows the height H(t), in feet, of a projectile launched vertically from s feet above the ground after t seconds. The initial speed of the projectile is v feet per second. Part A: The projectile was launched from a height of 96 feet with an initial velocity of 80 feet per second. Create an equation to find the time taken by the projectile to fall on the ground. (2 points) Part B: What is the maximum height that the projectile will reach? Show your work. (2 points)

OpenStudy (anonymous):

Part C: Another object moves in the air along the path of g(t) = 31 + 32.2t where g(t) is the height, in feet, of the object from the ground at time t seconds. Use a table to find the approximate solution to the equation H(t) = g(t), and explain what the solution represents in the context of the problem? [Use the function H(t) obtained in Part A, and estimate using integer values] (4 points) Part D: Do H(t) and g(t) intersect when the projectile is going up or down, and how do you know? (2 points)

OpenStudy (anonymous):

I dont know this but I will get my freands mk

OpenStudy (anonymous):

Awesome, thx so much! =)

OpenStudy (anonymous):

np

OpenStudy (anonymous):

@SyedMohammed98

OpenStudy (anonymous):

@k_lynn @Unanimoose

OpenStudy (anonymous):

@oOKawaiiOo

OpenStudy (anonymous):

@Bazer1hunndo

OpenStudy (anonymous):

I think Part A will be like this: \[h(t)= -16t^2 +96t +80\] Not sure =)

OpenStudy (anonymous):

Part A: h(t) = -16t^2 + vt + s h(t) = -16t^2 + (80)t + (96) h(t) = -16t^2 + 80t + 96 Does This help?:)

OpenStudy (anonymous):

Okay awesome, yes that helps =)

OpenStudy (anonymous):

What about Part B? @wolfd.andi

syed98 (syedmohammed98):

https://answers.yahoo.com/question/index?qid=20140407192241AAPFcvZ this might help :)

OpenStudy (anonymous):

Awesome thx @SyedMohammed98 =)

syed98 (syedmohammed98):

You're Welcome =)

OpenStudy (anonymous):

Part B, C, D anyone???? Please?!

OpenStudy (anonymous):

Nvr Mind =)